OpenStudy (anonymous):

What is the length of BC in the right triangle below? A. 142 B. 4222 C. 2640 D. 122 E. 14,844 F. 98

OpenStudy (anonymous):

picture please? :3

OpenStudy (anonymous):



OpenStudy (anonymous):

use pythagorean theoreom

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

you know how to work it out?

OpenStudy (anonymous):

No /.\ Do you mind helping me with steps .

OpenStudy (anonymous):

sure. let a=22 b=120 and we are looking c in this case so c=?

OpenStudy (anonymous):

so plug those value in the pythagorean theorem.A is the shortes side B is middle and C is the hypothenuse

OpenStudy (anonymous):

22^2+120^2=c^2

OpenStudy (anonymous):

484+14400=14884?

OpenStudy (anonymous):

mm yeah but remember that in the formula c have ^2 which means square root.So find the square root of 14884

OpenStudy (anonymous):

I got its 122! Thank you Josh (: @joshhfers